# Presenting a STEM Ways of Thinking Framework for Engineering Design-based Physics Problems

### v3

@@ 1 Introduction
Among the several innovations in educational practices is the approach of teaching STEM (science, technology, engineering, and mathematics) by integrating multiple disciplines [8, 9]. As educators and researchers, we have a responsibility towards preparing students in being “well informed about technical matters and well educated in the STEM subjects” [10, preface.vii]. Students need to be provided both opportunities and guidance to develop a variety of thinking skills. In this context, the notion of ‘Ways of Thinking’ (WoT) has gained traction in STEM education. A few researchers [11, 12, 13, 14] have even proposed WoT frameworks in the context of STEM-based problem solving and Engineering Education Research. Our goal in this paper is to propose our own WoT framework, designed specifically for Engineering Design (ED)-based physics learning.
Our confidence in presenting a fresh framework stems from a series of prior qualitative studies [15, 16] on the ‘design-science gap’ [17, 18, 19] in engineering design (ED)-based [20] projects in an introductory physics laboratory. Most notably, in our recently accepted manuscript [21], we had presented an elaborate analytic framework to explore, examine, and operationalize ‘design thinking’ and ‘science thinking’, drawing heavily from various qualitative research methodologies. We investigated how students engaged in design thinking, science thinking, mathematical thinking, and metacognitive reflection while solving an instructor-assigned ED problem. On the basis of our analysis, we suggested transitioning from the deficit-framed notion of design-science ‘gap’ to the asset-framed notion of design-science ‘connection’ (see Table 1).
@@
In line with the definitions proposed by Dalal et al. [11, p.109] and English [14, p.1220], we define ‘Ways of Thinking’ in the context of ED-based projects in physics as the approaches students adopt to intentionally think, act, and engage throughout their design projects, guiding their decision-making processes. These interconnected ways of thinking constitute a comprehensive framework for exploring, analyzing, evaluating, reflecting, and solving complex, interdisciplinary challenges with a focus on practical solutions. In our context, ‘Ways of Thinking’ illuminate how students approach engineering design problems, apply physics concepts, make decisions, and engage in iterative processes to develop conceptual solutions and deepen their understanding. Through this study, we aim to introduce a Ways of Thinking framework into Physics Education Research (PER), and in the process extend the discussion on development and significance of such frameworks.
While elaborating on the elements of our framework, we draw upon a range of reform documents with particular emphasis on physics education and PER. Based on our experience both as educators and researchers, we are confident our framework, in addition to extending the discourse on WoT frameworks, offers valuable insights for multiple stakeholders: educators, researchers, and students. WoT frameworks depend on educational contexts which, arguably, are inherently complex. It is essential that we remain open to, and learn from, a spectrum of perspectives. To this end, we engage in a detailed comparison with two recent contemporary WoT frameworks. In this study, we have two specific research goals (RGs).
@@ 2.1 The National Research Council report - 2013
Langenberg, the committee chair, was profuse in his appreciation of the members for their “deep dedication to physics and the ways of thinking that characterize it” [10, Preface, p.viii]. The report, while making recommendations to improve undergraduate physics education, goes on to urge educators “to actively engage students in the learning process, paying attention to their spontaneous ways of thinking…” [10, Chapter 3, p.75]. The lead author of the current study drew significant inspiration from the emphasis on the words “characterize” and “spontaneous”, and the repeated use of the phrase “ways of thinking” which appears at least seven times in the report. However, the report does not operationally define “ways of thinking”.
@@
Even though the NRC report [10], has not defined “ways of thinking”, several researchers have explored the notion of ‘Ways of Thinking’ in STEM and Engineering Education Research (EER) contexts. Denick and colleagues [23] identified ways of thinking such as science thinking, technology thinking, engineering thinking, and mathematical thinking. Acknowledging that identifying strict boundaries between these ways of thinking would be challenging, they recommend a holistic integration of these ways of thinking to promote student learning. Notably, in addition to discussing ‘science thinking’, they also reference ‘engineering design thinking’, both of which were explored in considerable depth in our recent studies [15, 22, 21].
It is our considered view that educators must feel empowered to interpret and apply the notion of ‘Ways of Thinking’ in ways that best resonate with their individual teaching styles, unique educational contexts and objectives, disciplinary practices, and the specific needs of their students. By adopting a flexible approach, educators can better support diverse learning outcomes and promote a deeper understanding of, and integration across, STEM disciplines.
@@ 2.4 WoT Frameworks - Recent Developments
In what appears to be a phase of significant development from 2021 to 2023, we identified three studies which saw a movement towards ‘characterizing’ WoT in contexts which have some commonalities with ours.
@@
In 2022, Slavit et al. [13, 24] presented an analytic framework for understanding students’ thinking in STEM environments, based on their previous studies on how students engage in claim, evidence, and reasoning. We were particularly impressed with their detailed descriptions on how they developed the coding schema for their studies and their justifications for coding decisions. They emphasize the potential of SWoT frameworks as valuable tools for educators, capable of capturing “transdisciplinary thinking” [24, p.145] and fostering greater integration across various disciplinary content areas. Most importantly, they indicate a scope for including ‘computational thinking’ [24, p.134] into SWoT frameworks.
@@ 2.5 Why a new SWoT Framework?
Figure 1: The proposed Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework. The double-headed arrows illustrate the interconnectedness of each element with the other four Ways of Thinking. A non-exhaustive list of pedagogical goals that the framework can support is presented. The concept of ‘Learning Innovation’ is adapted from English [14].
@@
In 2022 the lead author of the current study embarked on an exploratory study [15] into what we termed as the “Types of Thinking” that students engage in within an engineering design (ED)-based laboratory task. An ED problem [20] is client-driven and goal-oriented, set in an authentic context to engage students. It includes constraints such as cost, time, materials, familiar resources, and tools. Given that this is a physics course, we structured the problem to facilitate the application of physics concepts. The solution is a tangible product or process. Another important characteristic of ED problems is that they have multiple solution paths, requiring teamwork to creatively balance constraints and resources [25, p.60, 61].
@@
In 2023, we focused on characterizing the design-science gap in students’ thinking within an ED task [16]. Building on this study, in 2024, inspired by ongoing emphasis on SWoT, we transitioned to using the term ‘STEM Ways of Thinking’ [22] and examined how student groups’ thinking compared when solving both instructor-assigned and student-generated ED problems. Our recently accepted manuscript [21] presents an in-depth investigation of the design-science connections in students’ thinking while solving an instructor-assigned ED problem (see Figure 2) in a physics laboratory. To guide our analysis, we developed and applied a rigorous analytic framework, integrating qualitative research approaches such as traditional coding, thematic analysis, and thick description. Given the lack of universally accepted definitions of ‘design thinking’ and ‘science thinking’, a key objective was to characterize these constructs within our context. Having established a detailed characterization of students’ SWoT in ED-based tasks, we saw an opportunity to contribute to the broader discourse on SWoT frameworks, leading to the development of the WoT4EDP framework.
@@
At this juncture, we confronted a critical question: Is the development of a new framework truly necessary, or would modifying existing ones suffice? As we explain in subsequent sections, the frameworks proposed by Dalal et al. and English were originally designed for contexts that differ significantly from ours. The elements of Dalal et al.’s and English’s frameworks, in our view, do not always directly apply to physics classrooms, though valuable connections can be made depending on local contexts. Our aim is to propose a framework closely aligned with ED-based physics courses while maintaining adaptability for broader application in physics education. Furthermore, we sought to ground our framework in reform documents and research articles specific to Physics Education Research (PER), which may not have been a primary consideration of Dalal et al. and English. Notably, neither framework provides detailed, practical guidance on how to apply their models to actual student data, at least at the time of submission of this manuscript. Additionally, neither incorporates computational thinking—a key aspect in our pedagogical interventions. We also emphasize the interconnectedness of the five elements within our framework, a dimension that may not be as thoroughly explored in their works. While we acknowledge the influence of these existing frameworks, our approach extends beyond merely adding to them.
@@ 2.6 Guiding Perspectives
In addressing the question, “Why Are You Doing This Study?” [26, ch.2], we were guided by Maxwell’s Qualitative Research Design (2013). We strove earnestly to align our study with the eight key markers of quality in qualitative research outlined by Sarah Tracy: (a) worthy topic, (b) rich rigor, (c) sincerity, (d) credibility, (e) resonance, (f) significant contribution, (g) ethics, and (h) meaningful coherence [27, p.840]. To ensure ‘reliability’ and ‘validity’ [28, 27, 29, 30] of our framework development process, we anchored our approach in widely recognized reform documents published by organizations such as the National Research Council (NRC), the American Association of Physics Teachers (AAPT), the Mathematical Association of America (MAA), the College Board, and others, along with contemporary research articles. These sources provided a solid foundation for identifying and defining the five elements of our framework, shaping its conceptualization in a manner relevant to Physics Education and our specific educational context [26, ch.3].
Regarding the definitions of the five elements—design-based, science-based, mathematics-based, metacognitive reflection, and computational thinking—we recognize that there is no universally agreed-upon definition for such broad and complex concepts. In this context, we found reassurance in the philosophical view that “there is no possibility of attaining a single, ’correct’ understanding of the world” [31]. In her evocative essay The Role of the Researcher in Qualitative Psychology Kiegelmann argues that researcher’s perspective has an important influence on any qualitative study. She aptly quotes Putnam’s succinct words that there is no ultimate “God’s eye view” [31, p.15]. (Kiegelmann, 2002, p. 15). Accordingly, we advocate for a flexible approach that allows these concepts to be adapted to specific educational contexts. These elements are conceptual constructs rather than fixed, measurable entities, and their interpretation may vary depending on the learning environment and objectives. This inherent flexibility enhances the framework’s capacity to address the complexities of real-world educational settings.
@@
Consistent with our perspective that there can be no one-size-fits-all WoT framework - particularly in educational contexts - and that human understanding is enriched by having diverse viewpoints, we found Weiss’ approach of viewing “evaluation as enlightenment” [32, p.223] to be particularly valuable. In light of the complexities inherent in educational settings, we also resonate with Scriven’s assertion that “the most complete picture of reality can be constructed through the use of multiple, different perspectives” [32, p.222].
@@ 3 Presenting the WoT4EDP Framework
Transcripts of group discussions within 14 student groups in response to prompts. To explore how student-groups’ ways of thinking compare while engaging in two multi-week ED tasks- an ‘instructor-assigned’ ED problem and a ‘student-generated’ ED problem. [22]
@@
Transcripts of group discussions and final written reports of 14 student-groups To characterize student-groups’ ‘design thinking’ and ‘science thinking’. To explore the connections between student-groups’ design and science thinking. To investigate the influence of scaffolds on student-groups’ design and science thinking [21] (manuscript accepted for publication).
@@ 3.2 Engineering Design based Labs
In parallel with the hands-on and simulation-based lab tasks, student groups of two or three engaged with an ED problem. Seven essential characteristics of ED, as outlined by Capobianco et al. [25], guided the development of our ED problem: client-driven and goal-oriented, authentic context, constraints, use of familiar materials, resources, and tools, solution is an artifact or process, multiplicity of solutions, and teamwork. In the context of our study, we adopt the definition of ED as “a recursive activity that results in artifacts -- physical or virtual -- as well as processes” [20, p.348]. This includes stages such as: problem scoping and information gathering, idea generation, project realization, communication and documentation of performance results, and optimization. Similar terms such as “ill-structured” [40, p.444], “wicked” [41], and “ill-defined” [42, p.226] problems also find mention in the literature, and, for our purposes, we shall not delve into any subtle differences that may exist among them. One caveat is that our students did not develop an actual physical setup due to time and resource constraints - typical aspects which may be obvious to those who teach large enrollment courses. We, therefore, focused on how students engaged in the ED process to progress to the solution.
@@ 3.3 The Road to WoT4EDP
In our initial study [15] we identified four distinct ‘Types of Thinking’: design, science, mathematical, and metacognitive, using qualitative coding of student work. Building on this foundation, we [33] incorporated computational thinking, leveraging student data from a similar ED-based laboratory sequence to further refine the thinking framework. Further, [16], we investigated the ‘design–science gap’ observed in student projects, providing insight into how these modes of thinking interact. Inspired by ongoing discourse on thinking frameworks [11, 24, 13, 14], we adopted the term ‘Ways of Thinking’ in our next study [22], aligning our framework with contemporary educational research. These efforts culminated in an article (under review) on the ‘design–science connection’ [21], where an elaborate coding scheme was developed to characterize and contextualize student thinking in ED-based physics projects.
@@
In our studies, we utilized qualitative methods, including coding and thematic analysis, to examine student artifacts such as transcripts of group discussions and written reports (see Table 1). To analyze our data, we developed an inductive coding scheme guided by the Gioia framework [46], integrating thematic analysis [47] informed by the MIRACLE framework [48]. This approach enabled us to provide ‘thick descriptions’ [49, 50] of our findings.

### v4

@@ 1 Introduction
Among the several innovations in educational practices is the approach of teaching STEM (science, technology, engineering, and mathematics) by integrating multiple disciplines [8, 9]. As educators and researchers, we have a responsibility towards preparing students in being “well informed about technical matters and well educated in the STEM subjects” [10, preface.vii]. Students need to be provided both opportunities and guidance to develop a variety of thinking skills. In this context, the notion of Ways of Thinking (WoT) has gained traction in STEM education. A few researchers [11, 12, 13, 14] have even proposed WoT frameworks in the context of STEM-based problem solving and Engineering Education Research. Our goal in this paper is to propose our own WoT framework, designed specifically for Engineering Design (ED)-based physics learning.
Our confidence in presenting a fresh framework stems from a series of prior qualitative studies [15, 16] on the design-science gap [17, 18, 19] in engineering design (ED)-based [20] projects in an introductory physics laboratory. Most notably, in our recent paper [21], we had presented an elaborate analytic framework to explore, examine, and operationalize design thinking and science thinking, drawing heavily from various qualitative research methodologies. We investigated how students engaged in design thinking, science thinking, mathematical thinking, and metacognitive reflection while solving an instructor-assigned ED problem. On the basis of our analysis, we suggested transitioning from the deficit-framed notion of design-science gap to the asset-framed notion of design-science connection (see Table 1).
@@
In line with the definitions proposed by Dalal et al. [11, p.109] and English [14, p.1220], we define Ways of Thinking in the context of ED-based projects in physics as the approaches students adopt to intentionally think, act, and engage throughout their design projects, guiding their decision-making processes. These interconnected ways of thinking constitute a comprehensive framework for exploring, analyzing, evaluating, reflecting, and solving complex, interdisciplinary challenges with a focus on practical solutions. In our context, Ways of Thinking illuminate how students approach engineering design problems, apply physics concepts, make decisions, and engage in iterative processes to develop conceptual solutions and deepen their understanding. Through this study, we aim to introduce a Ways of Thinking framework into Physics Education Research (PER), and in the process extend the discussion on development and significance of such frameworks.
While elaborating on the elements of our framework, we draw upon a range of reform documents [22, 23, 24, 25, 26, 27] with particular emphasis on physics education and PER. Based on our experience both as educators and researchers, we are confident our framework, in addition to extending the discourse on WoT frameworks, offers valuable insights for multiple stakeholders: educators, researchers, and students. WoT frameworks depend on educational contexts which, arguably, are inherently complex. It is essential that we remain open to, and learn from, a spectrum of perspectives. To this end, we engage in a detailed comparison with two recent contemporary WoT frameworks. In this study, we have two specific research goals (RGs).
@@ 2.1 The National Research Council report - 2013
Langenberg, the committee chair, was profuse in his appreciation of the members for their “deep dedication to physics and the ways of thinking that characterize it” [10, Preface, p.viii]. The report, while making recommendations to improve undergraduate physics education, goes on to urge educators “to actively engage students in the learning process, paying attention to their spontaneous ways of thinking…” [10, Chapter 3, p.75]. The lead author of the current study drew significant inspiration from the emphasis on the words characterize and spontaneous, and the repeated use of the phrase ways of thinking which appears at least seven times in the report. However, the report does not operationally define ways of thinking.
@@
Even though the NRC report [10], has not defined ways of thinking, several researchers have explored the notion of Ways of Thinking in STEM and Engineering Education Research (EER) contexts. Denick and colleagues [29] identified ways of thinking such as science thinking, technology thinking, engineering thinking, and mathematical thinking. Acknowledging that identifying strict boundaries between these ways of thinking would be challenging, they recommend a holistic integration of these ways of thinking to promote student learning. Notably, in addition to discussing science thinking, they also reference engineering design thinking, both of which were explored in considerable depth in our recent studies [15, 28, 21].
It is our considered view that educators must feel empowered to interpret and apply the notion of Ways of Thinking in ways that best resonate with their individual teaching styles, unique educational contexts and objectives, disciplinary practices, and the specific needs of their students. By adopting a flexible approach, educators can better support diverse learning outcomes and promote a deeper understanding of, and integration across, STEM disciplines.
@@ 2.4 WoT Frameworks - Recent Developments
In what appears to be a phase of significant development from 2021 to 2023, we identified three studies which saw a movement towards characterizing WoT in contexts which have some commonalities with ours.
@@
In 2022, Slavit et al. [13, 30] presented an analytic framework for understanding students’ thinking in STEM environments, based on their previous studies on how students engage in claim, evidence, and reasoning. We were particularly impressed with their detailed descriptions on how they developed the coding schema for their studies and their justifications for coding decisions. They emphasize the potential of SWoT frameworks as valuable tools for educators, capable of capturing “transdisciplinary thinking” [30, p.145] and fostering greater integration across various disciplinary content areas. Most importantly, they indicate a scope for including computational thinking [30, p.134] into SWoT frameworks.
@@ 2.5 Why a new SWoT Framework?
Figure 1: The proposed Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework. The double-headed arrows illustrate the interconnectedness of each element with the other four Ways of Thinking. A non-exhaustive list of pedagogical goals that the framework can support is presented. The concept of Learning Innovation is adapted from English [14].
@@
In 2022 the lead author of the current study embarked on an exploratory study [15] into what we termed as the “Types of Thinking” that students engage in within an engineering design (ED)-based laboratory task. An ED problem [20] is client-driven and goal-oriented, set in an authentic context to engage students. It includes constraints such as cost, time, materials, familiar resources, and tools. Given that this is a physics course, we structured the problem to facilitate the application of physics concepts. The solution is a tangible product or process. Another important characteristic of ED problems is that they have multiple solution paths, requiring teamwork to creatively balance constraints and resources [31, p.60, 61].
@@
In 2023, we focused on characterizing the design-science gap in students’ thinking within an ED task [16]. Building on this study, in 2024, inspired by ongoing emphasis on SWoT, we transitioned to using the term STEM Ways of Thinking [28] and examined how student groups’ thinking compared when solving both instructor-assigned and student-generated ED problems. Our recent paper [21] presents an in-depth investigation of the design-science connections in students’ thinking while solving an instructor-assigned ED problem (see Figure 2) in a physics laboratory. To guide our analysis, we developed and applied a rigorous analytic framework, integrating qualitative research approaches such as traditional coding, thematic analysis, and thick description. Given the lack of universally accepted definitions of design thinking and science thinking, a key objective was to characterize these constructs within our context. Having established a detailed characterization of students’ SWoT in ED-based tasks, we saw an opportunity to contribute to the broader discourse on SWoT frameworks, leading to the development of the WoT4EDP framework.
@@
At this juncture, we confronted a critical question: Is the development of a new framework truly necessary, or would modifying existing ones suffice? As we explain in subsequent sections, the frameworks proposed by Dalal et al. and English were originally designed for contexts that differ significantly from ours. The elements of Dalal et al.’s and English’s frameworks, in our view, do not always directly apply to physics classrooms, though valuable connections can be made depending on local contexts. Our aim is to propose a framework closely aligned with physics courses that include ED-based experiences while maintaining adaptability for broader application in physics education. Furthermore, we sought to ground our framework in reform documents [22, 23, 24, 25, 26, 27] and research articles specific to Physics Education Research (PER), which may not have been a primary consideration of Dalal et al. and English. Notably, neither Dalal et al. nor English provide detailed, practical guidance on how to apply their models to actual student data, at least at the time of submission of this manuscript. Additionally, neither incorporates computational thinking—a key aspect in our pedagogical interventions. We also emphasize the interconnectedness of the five elements within our framework, a dimension that may not be as thoroughly explored in their works. While we acknowledge the influence of these existing frameworks, our approach extends beyond merely extending them.
@@ 2.6 Guiding Perspectives
In addressing the question, “Why Are You Doing This Study?” [32, ch.2], we were guided by Maxwell’s Qualitative Research Design (2013). We strove earnestly to align our study with the eight key markers of quality in qualitative research outlined by Tracy: (a) worthy topic, (b) rich rigor, (c) sincerity, (d) credibility, (e) resonance, (f) significant contribution, (g) ethics, and (h) meaningful coherence [33, p.840]. To ensure reliability and validity (some researchers suggest the use of equivalent terms such as trustworthiness and dependability[34, 33, 35, 36] of our framework development process, we anchored our approach in widely recognized reform documents published by organizations such as the National Research Council (NRC), the American Association of Physics Teachers (AAPT), the Mathematical Association of America (MAA), the College Board, and others, along with contemporary research articles. These sources provided a solid foundation for identifying and defining the five elements of our framework, shaping its conceptualization in a manner relevant to Physics Education and our specific educational context [32, ch.3].
Regarding the definitions of the five elements—design-based, science-based, mathematics-based, metacognitive reflection, and computational thinking—we recognize that there is no universally agreed-upon definition for such broad and complex concepts. In this context, we found reassurance in the philosophical view that “there is no possibility of attaining a single, ’correct’ understanding of the world” [37]. In her evocative essay The Role of the Researcher in Qualitative Psychology Kiegelmann argues that researcher’s perspective has an important influence on any qualitative study. She aptly quotes Putnam’s succinct words that there is no ultimate “God’s eye view” [37, p.15]. (Kiegelmann, 2002, p. 15). Accordingly, we advocate for a flexible approach that allows these concepts to be adapted to specific educational contexts. These elements are conceptual constructs rather than fixed, measurable entities, and their interpretation may vary depending on the learning environment and objectives. This inherent flexibility enhances the framework’s capacity to address the complexities of real-world educational settings.
@@
Consistent with our perspective that there can be no one-size-fits-all WoT framework - particularly in educational contexts - and that human understanding is enriched by having diverse viewpoints, we found Weiss’ approach of viewing “evaluation as enlightenment” [38, p.223] to be particularly valuable. In light of the complexities inherent in educational settings, we also resonate with Scriven’s assertion that “the most complete picture of reality can be constructed through the use of multiple, different perspectives” [38, p.222].
@@ 3 Presenting the WoT4EDP Framework
Transcripts of group discussions within 14 student groups in response to prompts. To explore how student-groups’ ways of thinking compare while engaging in two multi-week ED tasks- an instructor-assigned ED problem and a student-generated ED problem. [28]
@@
Transcripts of group discussions and final written reports of 14 student-groups To characterize student-groups’ design thinking and science thinking. To explore the connections between student-groups’ design and science thinking. To investigate the influence of scaffolds on student-groups’ design and science thinking [21].
@@ 3.2 Engineering Design based Labs
In parallel with the hands-on and simulation-based lab tasks, student groups of two or three engaged with an ED problem. Seven essential characteristics of ED, as outlined by Capobianco et al. [31], guided the development of our ED problem: client-driven and goal-oriented, authentic context, constraints, use of familiar materials, resources, and tools, solution is an artifact or process, multiplicity of solutions, and teamwork. In the context of our study, we adopt the definition of ED as “a recursive activity that results in artifacts -- physical or virtual -- as well as processes” [20, p.348]. This includes stages such as: problem scoping and information gathering, idea generation, project realization, communication and documentation of performance results, and optimization. Similar terms such as “ill-structured” [46, p.444], “wicked” [47], and “ill-defined” [48, p.226] problems also find mention in the literature, and, for our purposes, we shall not delve into any subtle differences that may exist among them. One caveat is that our students did not develop an actual physical setup due to time and resource constraints - typical aspects which may be obvious to those who teach large enrollment courses. We, therefore, focused on how students engaged in the ED process to progress to the solution.
@@ 3.3 The Road to WoT4EDP
In our initial study [15] we identified four distinct Types of Thinking: design, science, mathematical, and metacognitive, using qualitative coding of student work. Building on this foundation, we [39] incorporated computational thinking, leveraging student data from a similar ED-based laboratory sequence to further refine the thinking framework. Further, [16], we investigated the design–science gap observed in student projects, providing insight into how these modes of thinking interact. Inspired by ongoing discourse on thinking frameworks [11, 30, 13, 14], we adopted the term Ways of Thinking in our next study [28], aligning our framework with contemporary educational research. These efforts culminated in our recent paper on design–science connection [21], where an elaborate coding scheme was developed to characterize and contextualize student thinking in ED-based physics projects.
@@
In our studies, we utilized qualitative methods, including coding and thematic analysis, to examine student artifacts such as transcripts of group discussions and written reports (see Table 1). To analyze our data, we developed an inductive coding scheme guided by the Gioia framework [50], integrating thematic analysis [51] informed by the MIRACLE framework [52]. This approach enabled us to provide thick descriptions [53, 54] of our findings.